OpenStudy (anonymous):

Which is (sin B+ cos C) in the right triangle ABC, given AB = 20, AC = 48, and mA = 90? and the answer has to be a fraction

OpenStudy (id21):

Find hypotenuse first \[BC=\sqrt{20^2+48^2}=52\] sinB=AC/BC=48/52 cosC=AC/BC=48/52 sinB+cosC=48/52+48/52=24/13